Evaluate the following

Log381

Log273

Log644

Log901

1Log


1
Expert's answer
2020-09-28T19:28:13-0400

There are some useful properties of logarithms that we will use: "\log_{a^k}a = \frac{1}{k}" , "\log_{a}a^m = m", "\log_a1=0".


"\displaystyle \log_381=\log_3 3^4=4"

"\displaystyle \log_{27}3 = \log_{3^3}3=\frac{1}{3}"

"\displaystyle \log_{64} 4 = \log_{4^3}4= \frac{1}{3}"

"\log_{90}1=0"